OpenStudy (anonymous):

(3v+2)^2 find the product

OpenStudy (alfie):

9v^2 + 12v + 4

OpenStudy (anonymous):

use the expansion formula\[(a+b)^2=a^2+2ab+b^2\]in your problem a=3v and b=2; that's how alfie got the correct answer

OpenStudy (anonymous):

so is it like 3v+2)(3v+2)

OpenStudy (anonymous):

yes; using the formula is just easier with a little practice

OpenStudy (anonymous):

thanks